A coin placed on a turntable rotating at 33.0 rev/min will stay there if its center is placed no further than 8.8 cm from the axis of rotation. What is the maximum distance the center of the coin can be placed from the axis if the turntable rotates at 46.0 rev/min? What is the coefficient of friction between the coin and turntable?
Answer #1

Centripetal force = m*a

mue_s*N = m*r*w^2

mue_s*m*g = m*r*w^2

mue_s = r*w^2/g

in both cases static friction is same

so,

r2*w2^2 = r1*w1^2

r2 = r1*(w1/w2)^2

= 8.8*(33/46)^2

= 4.53 cm